# Photometry and Spectroscopy of SN 2024pxl: A Luminosity Link Among Type Iax Supernovae

### v1

@@ 2.2.2 NIR Spectra
We also obtained two NIR spectra of SN 2024pxl with the Espectrógrafo Multiobjeto Infra-Rojo spectrograph (EMIR; Garzón et al. 2022) mounted on the GTC, which were reduced using a dedicated pipeline based on PyEMIR (Pascual et al., 2010; Cardiel et al., 2019). The Southern Astrophysical Research (SOAR) telescope, equipped with TripleSpec (Schlawin et al., 2014), was also deployed to obtain NIR spectra. Spectra acquired by SOAR were reduced using the IDL-based Spextool package (Cushing et al., 2004). All NIR observations for SN 2024pxl are reported in Table 7.
3 Distance, extinction, and explosion epoch
SN 2024pxl is hosted by NGC 6384 at a spectroscopic redshift of (Lagattuta et al., 2013). Fortunately, a normal (though highly dust-extinguished) SN Ia with a well-sampled light curve from LCO and Lick Observatory, SN 2017drh, was also hosted by NGC 6384 (Stahl et al., 2019). We estimate the distance to SN 2017drh using BayeSN, a hierarchical Bayesian SN Ia light-curve model (Mandel et al., 2022; Grayling et al., 2024). SN 2017drh was located near the center of NGC 6384 and is significantly dust reddened, much more than SNe Ia typically used for cosmological analyses. Thus, we favored using BayeSN because it explicitly models host-galaxy dust reddening and extinction separately from intrinsic SN spectral energy distribution (SED) variations, unlike alternatives like SALT2 (Guy et al., 2007) that aggregate intrinsic and extrinsic SN Ia color variations.
@@ 3 Distance, extinction, and explosion epoch
The extinction experienced by SN 2024pxl due to the Milky Way in the direction of NGC 6384 is mag, corresponding to mag (Schlafly & Finkbeiner, 2011) and assuming = 3.1. Estimating the host-galaxy extinction is more difficult because, unlike the normal SN Ia 2017drh, we do not have a standard model color for the diverse Type Iax class. However, in our spectroscopic observations of SN 2024pxl with sufficient spectral resolution, we note the presence of narrow interstellar Na I D absorption lines from both the Milky Way and NGC 6384, with approximately equal strength. Thus, we crudely infer the host-galaxy contribution to the extinction of SN 2024pxl to be similar to the Milky Way extinction and adopt a total reddening and extinction along the line of sight to SN 2024pxl of mag and mag.
@@
indicating a shallower rise than the canonical case.
4 Light-Curve Properties
4.1 Light Curves and Color Curves
Figure 2 presents the evolution of the light curve for SN 2024pxl in the BgVriz bands, with dense sampling around maximum brightness in all bands. We use a low-order polynomial fit to estimate the time and magnitude at maximum light in the BgVriz bands. Additionally, we estimate the decline in magnitude from the light-curve peak to 15 days after () for these bands. These light-curve parameters for SN 2024pxl are listed in Table 1. Figure 3 displays the evolution of SN 2024pxl in the Swift UVOT bands. All bands shown in Figure 3, except for UVW2 and UVM2, cover the peak well. We estimate the Swift UVOT light-curve data parameters in Table 2.
| SN 2024pxl | band | band | band | band | band | band |
| JD of maximum light (2,460,000+) | 524.60.5 | 525.20.5 | 527.40.5 | 528.140.5 | 530.50.5 | 531.70.5 |
| Magnitude at maximum (mag) | 16.420.02 | 16.120.01 | 15.910.02 | 15.580.01 | 15.500.01 | 15.470.07 |
| Absolute magnitude at maximum (mag) | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.21 |
| (mag) | 1.790.09 | 1.510.04 | 0.840.06 | 0.480.04 | 0.530.04 | 0.490.10 |
@@ 4.1 Light Curves and Color Curves
4.2 Bolometric Light Curve
The pseudobolometric BgVriz light curve of SN 2024pxl is generated using SuperBol (Nicholl, 2018), incorporating the distance and extinction values outlined in Section 3. SuperBol takes the dereddened fluxes and uncertainties across passbands to construct an SED at each epoch. Luminosities are determined by integrating this spectrum while propagating uncertainties. To account for missing passbands in the UV and NIR, SuperBol fits a blackbody to each SED to derive a correction which is then applied to the pseudobolometric luminosities.
The pseudobolometric light curves of the other comparison SNe presented in Figure 7 are constructed using a similar approach. SN 2018cni (Singh et al., 2023), another bright Type Iax, is also included for comparison in Figure 7. The peak pseudobolometric BgVriz luminosity of SN 2024pxl is erg s-1. SN 2024pxl is less luminous than all the high-luminosity Type Iax SNe, such as SNe 2012Z, 2020rea, and 2020udy, but is brighter than the intermediate-luminosity Type Iax SN 2019muj. SN 2024pxl lies in the gap between the high-luminosity and low-luminosity Type Iax SNe, in agreement with our decline-rate comparison (Section 4.1).
We conduct analytical modeling of the pseudobolometric light curve for SN 2024pxl using the methods outlined by Arnett (1982) and Valenti et al. (2008). The basic assumptions of this model consist of a small initial radius, constant optical opacity, spherically symmetric and optically thick ejecta, and the inclusion of 56Ni in the ejected material. By fitting the pseudobolometric light curve of SN 2024pxl, we estimate a 56Ni mass of M⊙, an ejecta mass of M⊙, and a rise time of 10 days, with uncertainties derived from the covariance matrix. For the fit, we assume a constant opacity of cm2 g-1 and a photospheric velocity of 5000 km s-1 at maximum brightness. Directly estimating the photospheric expansion velocity from the bottom panel of Figure 8, we calculate km s-1 over the nearly linear expansion period from maximum light to 15 days. It is surprising that this value is larger than the typical photospheric absorption-line velocities, perhaps suggesting that the optical depth approaches unity farther out in the ejecta than the lines would predict.
@@ 4.2 Bolometric Light Curve
For Type Iax SNe, contributions from the unobserved UV and IR passbands remain poorly constrained. However, based on available data for a few Type Iax SNe, the combined UV and NIR contributions to the total bolometric flux have been estimated. Phillips et al. (2007) demonstrated that the UV flux accounts for approximately 20% of the UV-through-IR (UVOIR) light curve for an early photospheric phase in the case of SN 2005hk. Similarly, Yamanaka et al. (2015) estimated that the NIR band makes up about 20% of the optical+NIR flux around peak light for SN 2012Z. Tomasella et al. (2016) found that the UV+IR flux collectively accounted for % of the total emission in SN 2014ck. Moreover, Srivastav et al. (2020) and Dutta et al. (2022) calculated the ratio of (peak pseudobolometric luminosity) to (peak blackbody luminosity) as 0.69 for SN 2019gsc and 0.62 for SN 2020sck, respectively. In the case of SN 2020kyg, Srivastav et al. (2022) reported that about 60% of the total bolometric luminosity near maximum light is attributed to the optical emission.
@@
We have Swift observations that capture the light-curve evolution of SN 2024pxl around the peak. Using SuperBol for , , , and filters along with BgVriz filters, we find that the ratio of peak pseudobolometric luminosity using BgVriz and SDAUBgVriz bands (, , , ) is , suggesting that the UV may be contributing up to % of the bolometric luminosity near peak. Using the same analytical model for the SDAUBgVriz pseudobolometric light curve, we find that M⊙ of 56Ni is synthesized during the explosion of SN 2024pxl. This is in agreement with our earlier estimate using optical data only. Based on the observed UV+optical data for SN 2024pxl, our final estimate for the 56Ni mass is M⊙.
@@
Fink et al. (2014) presented numerical calculations for deflagration models of CO WDs with varying explosion energy (parametrized by the number of ignition kernels). These models span a broad range of explosion parameters (56Ni mass 0.03–0.38 M⊙, rise time 7.6–14.4 days). The estimated mass of 56Ni (0.03 M⊙) and rise time (10 days) for SN 2024pxl are consistent with the range provided by these models. However, the peak magnitude in the V band for SN 2024pxl ( mag) is on the fainter end of the estimation (16.84 to 18.96 mag) given by Fink et al. (2014). The mass of 56Ni in the case of SN 2024pxl matches with the 56Ni synthesized in the weakest-energy N1def (single ignition point) model presented by Fink et al. (2014). Radiative-transfer modeling based on the N1def model by Kwok et al. (2025) shows good agreement with the spectra of SN 2024pxl from the optical through the MIR. For low-luminosity Type Iax SNe, Kromer et al. (2015) studied the deflagration of a hybrid CONe WD. Our estimated peak magnitude of SN 2024pxl in the B band ( mag) and mass of 56Ni (0.03 M⊙) exceed the predictions from the hybrid CONe WD model ( to mag, mass of 56Ni M⊙) that may apply to lower-luminosity objects.
5 Spectral Evolution
5.1 Spectral Features and Comparisons
@@
Figures 9, 10, and 11 present the spectral evolution (LABEL:tab:spectroscopic_observations_24pxl) of SN 2024pxl from approximately 8 to 89 days relative to Bmax. Spectra were obtained nearly daily for almost two months after maximum, producing a rich data set allowing detailed tracking of spectral evolution. We compare the spectral features of SN 2024pxl with those of other well-studied Type Iax SNe at similar pre-maximum, near-maximum, post-maximum, and late phases in Figures 12, 13, 14, and 15, respectively.
5.1.1 Pre-maximum spectra
@@
We observe that at early, prepeak times, SN 2024pxl has a flat continuum consistent with the low temperatures we infer from the photometry (see Figure 8). The early-time spectra of SN 2024pxl display features such as Fe III, Fe II, and Si II (Figure 12). In agreement with our findings that SN 2024pxl is a photometrically intermediate Type Iax SN, it is spectroscopically intermediate as well, sharing similarities with both high-luminosity Type Iax SNe (e.g., SNe 2012Z, 2005hk, and 2020udy) and low-luminosity objects (e.g., SNe 2010ae and 2008ha). Specifically, SN 2024pxl has weak Si II features but also exhibits a C II line at 6580 Å. In high-luminosity Type Iax SNe, the Si II line is relatively weak, and C II is generally not detected. In contrast, in low-luminosity Type Iax SNe these features are much more prominent. The Ca II NIR triplet is present in SN 2024pxl but is weaker than in SNe 2010ae and 2008ha. This combination of weak Si II and presence of C II suggests that SN 2024pxl also serves as a spectroscopic transitional link between high-luminosity and low-luminosity Type Iax SNe.
5.1.2 Near-maximum spectra
@@
5.1.3 Post-maximum and late-time spectra
Figure 14 shows the post-maximum spectra of SN 2024pxl in the context of other objects. At these epochs, SN 2024pxl shows remarkable similarity with SN 2019muj, an intermediate-luminosity SN Iax. During the post-maximum phase, the Si II and C II lines vanish, giving way to Fe and Co lines in the 5000-–7000 Å range. Iron-group elements (IGEs) and a strong Ca II NIR triplet dominate the post-maximum spectra, similar to other high- and intermediate-luminosity objects. Low-luminosity Type Iax SNe (e.g., SNe 2010ae and 2008ha, shown in Figure 14) evolve more rapidly and exhibit narrower spectral signatures than SN 2024pxl. The Cr II feature near 4800 Å and the Co II lines near 6500 Å in SN 2024pxl share strong resemblance with SNe 2002cx and 2019muj. Additionally, the Ca II NIR triplet in SN 2019muj is strikingly similar to that in SN 2024pxl.
We compare the late-time spectrum of SN 2024pxl obtained at 89 days with spectra of SNe 2008ge (Foley et al., 2010), 2014dt (Singh et al., 2018), and 2020udy at similar epochs (Figure 15). At late times, the spectral lines become narrower and several forbidden emission lines such as [Fe II], [Ni II], and [Ca II] emerge along with narrow permitted lines such as Fe II. The strength of these forbidden lines increases with the cooling of the ejecta. Owing to their intrinsically low luminosity, late-time spectra of Type Iax SNe are rare. However, the late-time spectra of Type Iax SNe vary widely, with some showing broad emission lines and others dominated by narrow features (Foley et al., 2016). Of the comparison SNe in Figure 15, SN 2024pxl most closely resembles SN 2014dt, exhibiting narrower lines than the higher-luminosity objects SNe 2020udy and 2008ge. The narrow features seen in the late-phase spectra of Type Iax SNe are thought to originate from the innermost regions (Maeda & Kawabata, 2022; Camacho-Neves et al., 2023). Additional late-time observations of Type Iax SNe across the luminosity distribution will help to understand the origin and nature of these spectral features more precisely.

### v2

@@ II.2.2 NIR Spectra
We obtained two NIR spectra of SN 2024pxl with the Espectrógrafo Multiobjeto Infra-Rojo spectrograph (EMIR; Garzón et al. 2022) mounted on the GTC, which were reduced using a dedicated pipeline based on PyEMIR (Pascual et al., 2010; Cardiel et al., 2019). The Southern Astrophysical Research (SOAR) telescope, equipped with TripleSpec (Schlawin et al., 2014), was also deployed to obtain NIR spectra. Spectra acquired by SOAR were reduced using the IDL-based Spextool package (Cushing et al., 2004). All NIR observations for SN 2024pxl are reported in Table 7.
III Distance, extinction, and explosion epoch
SN 2024pxl is hosted by NGC 6384 at a spectroscopic redshift of (Lagattuta et al., 2013). Fortunately, a normal (though highly dust-extinguished) SN Ia with a well-sampled light curve from LCO and Lick Observatory, SN 2017drh, was also hosted by NGC 6384 (Stahl et al., 2019b). We estimate the distance to SN 2017drh using BayeSN, a hierarchical Bayesian SN Ia light-curve model (Mandel et al., 2022; Grayling et al., 2024). SN 2017drh was located near the center of NGC 6384 and is significantly dust reddened, much more than SNe Ia typically used for cosmological analyses. Thus, we favored using BayeSN because it explicitly models host-galaxy dust reddening and extinction separately from intrinsic SN spectral energy distribution (SED) variations, unlike alternatives like SALT2 (Guy et al., 2007) that aggregate intrinsic and extrinsic SN Ia color variations.
@@ III Distance, extinction, and explosion epoch
We also perform a BayeSN fit to SN 2017drh allowing to vary, given that previous observations of heavily dust-extinguished SNe Ia sometimes indicate low (Krisciunas et al., 2006; Elias-Rosa et al., 2006, 2008; Wang et al., 2008; Misra et al., 2008; Amanullah et al., 2015), possibly due to scattering effects in the surrounding medium or an atypical dust grain size distribution (Wang, 2005; Goobar, 2008; Kawabata et al., 2014; Gao et al., 2015; Bulla et al., 2018). In that case we can recover a model fit for SN 2017drh with and mag that is nearly indistinguishable from the model, but with almost 1 mag less total extinction and thus a larger distance modulus, mag. This would imply a distance of Mpc, significantly farther than our adopted estimate. Unfortunately, owing to the lack of NIR data for SN 2017drh, we cannot break the degeneracy between and . We thus proceed with our distance estimate using but caution that the lower- fit and increased distance would imply a factor of 2 change in the derived luminosity of SN 2024pxl. We note that these distances bracket the distance to SN 2017drh derived by Hoogendam et al. (2025) of Mpc using the SNooPy code (Burns et al., 2011).
The extinction experienced by SN 2024pxl due to the Milky Way in the direction of NGC 6384 is mag, corresponding to mag (Schlafly & Finkbeiner, 2011) and assuming = 3.1. Estimating the host-galaxy extinction is more difficult because, unlike the normal SN Ia 2017drh, we do not have a standard model color for the diverse Type Iax class. However, in our spectroscopic observations of SN 2024pxl, we note the presence of narrow interstellar Na I D absorption lines from both the Milky Way and NGC 6384, with approximately equal strength. Thus, we crudely infer the host-galaxy contribution to the extinction of SN 2024pxl to be similar to the Milky Way extinction and adopt a total reddening and extinction along the line of sight to SN 2024pxl of mag and mag.
@@
indicating a shallower rise than the typical case. We emphasize that the quoted 0.10 day represents the formal statistical uncertainty from the power-law fit. In practice, the first detection occurs 0.96 days after this fitted zero point, underscoring that the true uncertainty is significantly larger, being influenced by systematic effects such as deviations from the assumed flux rise behavior, and possible undetected flux below the survey limits. Thus, while the formal error quantifies the fit precision, the absolute accuracy of the explosion epoch is less certain. This distinction does not affect the broader discussion in this work.
IV Light-Curve Properties
IV.1 Light Curves and Color Curves
Figure 2 presents the evolution of the light curve for SN 2024pxl in the uBgVrRiIz and clear bands, with dense sampling around maximum brightness in all bands. We use a low-order polynomial fit to estimate the time and magnitude at maximum light in the uBgVrRiIz bands. Additionally, we estimate the decline in magnitude from the light-curve peak to 15 days after () for these bands. These light-curve parameters for SN 2024pxl are listed in Table 1. Figure 3 displays the evolution of SN 2024pxl in the Swift UVOT bands. All bands shown in Figure 3, except for UVW2 and UVM2, cover the peak well. We estimate the Swift UVOT light-curve data parameters in Table 2.
| SN 2024pxl | band | band | band | band | band | band | band | band | band |
| JD of maximum light (2,460,000+) | 523.000.5 | 524.60.5 | 525.20.5 | 527.40.5 | 528.140.5 | 529.40.5 | 530.20.5 | 531.10.5 | 531.70.5 |
| Magnitude at maximum (mag) | 17.020.01 | 16.420.02 | 16.100.01 | 15.900.01 | 15.560.01 | 15.440.01 | 15.490.01 | 15.120.02 | 15.470.07 |
| Absolute magnitude at maximum (mag) | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.21 |
| (mag) | 2.540.05 | 1.800.07 | 1.560.04 | 0.870.06 | 0.490.05 | 0.660.03 | 0.510.02 | 0.400.04 | 0.460.08 |
@@ IV.1 Light Curves and Color Curves
IV.2 Bolometric Light Curve
The integrated SED light curve of SN 2024pxl is generated from the BgVriz photometry using SuperBol (Nicholl, 2018), incorporating the distance and extinction values outlined in Section III. SuperBol takes the dereddened fluxes and uncertainties across passbands to construct an SED at each epoch. Luminosities are determined by integrating this spectrum while propagating uncertainties. To account for missing passbands in the UV and NIR, SuperBol fits a blackbody to each SED to derive a correction which is then applied to the integrated SED luminosities.
The integrated SED light curve of the other comparison SNe presented in Figure 7 are constructed using a similar approach. SN 2018cni (Singh et al., 2023), another high-luminosity SN Iax, is also included for comparison in Figure 7. The peak integrated SED BgVriz luminosity of SN 2024pxl is erg s-1. SN 2024pxl is less luminous than all the high-luminosity Type Iax SNe, such as SNe 2012Z, 2020rea, and 2020udy, but is brighter than the intermediate-luminosity Type Iax SN 2019muj. SN 2024pxl lies in the gap between the high-luminosity and low-luminosity Type Iax SNe, in agreement with our decline-rate comparison (Section IV.1).
We conduct analytical modeling of the integrated SED light curve for SN 2024pxl using the methods outlined by Arnett (1982) and Valenti et al. (2008). The basic assumptions of this model consist of a small initial radius, constant optical opacity, spherically symmetric and optically thick ejecta, and the inclusion of 56Ni in the ejected material. By fitting the integrated SED light curve of SN 2024pxl, we estimate a 56Ni mass of M⊙, an ejecta mass of M⊙, and a rise time of 11 days, with uncertainties derived from the covariance matrix. We note that this small formal error represents only the statistical uncertainty from the fit and does not capture systematic effects arising from model assumptions. The true uncertainty is expected to be larger, and the value should be regarded as a model-dependent estimate.
For the fit, we assume a constant opacity of cm2 g-1 and a photospheric velocity of 5000 km s-1 at maximum brightness. Directly estimating the photospheric expansion velocity from the bottom panel of Figure 8, we calculate km s-1 over the nearly linear expansion period from maximum light to 15 days. It is surprising that this value is larger than the typical photospheric absorption-line velocities, perhaps suggesting that the optical depth approaches unity farther out in the ejecta than the lines would predict.
@@ IV.2 Bolometric Light Curve
For Type Iax SNe, contributions from the unobserved UV and IR passbands remain poorly constrained. However, based on available data for a few Type Iax SNe, the combined UV and NIR contributions to the total bolometric flux have been estimated. Phillips et al. (2007) demonstrated that the UV flux accounts for approximately 20% of the UV-through-IR (UVOIR) light curve for an early photospheric phase in the case of SN 2005hk. Similarly, Yamanaka et al. (2015) estimated that the NIR band makes up about 20% of the optical+NIR flux around peak light for SN 2012Z. Tomasella et al. (2016) found that the UV+IR flux collectively accounted for % of the total emission in SN 2014ck. Moreover, Srivastav et al. (2020) and Dutta et al. (2022) calculated the ratio of (peak integrated SED luminosity) to (peak blackbody luminosity) as 0.69 for SN 2019gsc and 0.62 for SN 2020sck, respectively. In the case of SN 2020kyg, Srivastav et al. (2022) reported that about 60% of the total bolometric luminosity near maximum light is attributed to the optical emission.
@@
We have Swift observations that capture the light-curve evolution of SN 2024pxl around the peak. Using SuperBol for , , , and filters along with BgVriz filters, we find that the ratio of peak integrated SED luminosity using BgVriz and SDAUBgVriz bands (, , , ) is , suggesting that the UV may be contributing up to % of the bolometric luminosity near peak. Using the same analytical model for the SDAUBgVriz integrated SED light curve, we find that M⊙ of 56Ni is synthesized during the explosion of SN 2024pxl. While this result is consistent with our earlier optical-only 56Ni mass estimate (0.026 0.002 M⊙), the quoted uncertainties represent formal errors from the fitting procedure and do not fully capture systematic effects associated with the simplifying assumptions of the analytic model (e.g., constant opacity, spherical symmetry).
@@
Fink et al. (2014) presented numerical calculations for deflagration models of CO WDs with varying explosion energy (parametrized by the number of ignition kernels). These models span a broad range of explosion parameters (56Ni mass 0.03–0.38 M⊙, rise time 7.6–14.4 days). The estimated mass of 56Ni (0.03 M⊙) and rise time (11 days) for SN 2024pxl are consistent with the range provided by these models. However, the peak magnitude in the V band for SN 2024pxl ( mag) is on the fainter end of the estimation (16.84 to 18.96 mag) given by Fink et al. (2014). The mass of 56Ni in the case of SN 2024pxl matches with the 56Ni synthesized in the weakest-energy N1def (single ignition point) model presented by Fink et al. (2014). Radiative-transfer modeling based on the N1def model by Kwok et al. (2025) shows good agreement with the spectra of SN 2024pxl from the optical through the MIR.
For low-luminosity Type Iax SNe, Kromer et al. (2015) studied the deflagration of a hybrid CONe WD. Our estimated peak magnitude of SN 2024pxl in the B band ( mag) and mass of 56Ni (0.03 M⊙) exceed the predictions from the hybrid CONe WD model ( to mag, mass of 56Ni M⊙) that may apply to lower-luminosity objects. These results also favor the interpretation that SN 2024pxl is an interesting luminosity link, with explosion parameters higher than those predicted for low-luminosity objects but lying at the lower end of the models proposed to explain the high-luminosity members of the class.
V Spectral Evolution
V.1 Spectral Features and Comparisons
@@
Figures 9, 10, and 11 present the spectral evolution (LABEL:tab:spectroscopic_observations_24pxl) of SN 2024pxl from approximately 9 to 89 days relative to Bmax. Spectra were obtained nearly daily for almost two months after maximum, producing a rich data set allowing detailed tracking of spectral evolution. We compare the spectral features of SN 2024pxl with those of other well-studied Type Iax SNe at similar pre-maximum, near-maximum, post-maximum, and late phases in Figures 12, 13, 14, and 15, respectively.
V.1.1 Pre-maximum spectra
@@
We observe that at early, prepeak times, SN 2024pxl have a flat continuum consistent with the low temperatures we infer from the photometry (see Figure 8). The early-time spectra of SN 2024pxl display features such as Fe III, Fe II, and Si II (Figure 12). In agreement with our findings that SN 2024pxl is a photometrically intermediate Type Iax SN, it is spectroscopically intermediate as well, sharing similarities with both high-luminosity Type Iax SNe (e.g., SNe 2012Z, 2005hk, and 2020udy) and low-luminosity objects (e.g., SNe 2010ae and 2008ha). Specifically, SN 2024pxl has weak Si II features but also exhibits a C II line at 6580 Å. In high-luminosity Type Iax SNe, the Si II line is relatively weak, and C II is generally not detected. In contrast, in low-luminosity Type Iax SNe, these features are much more prominent. The Ca II NIR triplet is present in SN 2024pxl but is weaker than in SNe 2010ae and 2008ha. This combination of weak Si II and presence of C II suggests that SN 2024pxl also serves as a spectroscopic transitional link between high-luminosity and low-luminosity Type Iax SNe.
V.1.2 Near-maximum spectra
@@
V.1.3 Post-maximum and late-time spectra
Figure 14 shows the post-maximum spectra of SN 2024pxl in the context of other objects. At these epochs, SN 2024pxl closely resembles SN 2019muj, an intermediate-luminosity Type Iax SN. During the post-maximum phase, the Si II and C II lines vanish, giving way to Fe and Co lines in the 5000-–7000 Å range. Iron-group elements (IGEs) and a strong Ca II NIR triplet dominate the post-maximum spectra, similar to other high- and intermediate-luminosity objects. Low-luminosity Type Iax SNe (e.g., SNe 2010ae and 2008ha, shown in Figure 14) evolve more rapidly and exhibit narrower spectral signatures than SN 2024pxl. The Cr II feature near 4800 Å and the Co II lines near 6500 Å in SN 2024pxl share resemblance with high-luminosity SNe 2002cx and 2019muj. Additionally, the Ca II NIR triplet in SN 2019muj is strikingly similar to that in SN 2024pxl.
We compare the late-time spectrum of SN 2024pxl obtained at 89 days with spectra of SNe 2008ge (Foley et al., 2010), 2014dt (Singh et al., 2018), and 2020udy at similar epochs (Figure 15). At late times, the spectral lines become narrower and several forbidden emission lines such as [Fe II], [Ni II], and [Ca II] emerge along with narrow permitted lines such as Fe II. The strength of these forbidden lines increases with the cooling of the ejecta (Foley et al., 2016). Owing to their intrinsically low luminosity, late-time spectra of Type Iax SNe are rare. The late-time spectra of Type Iax SNe vary widely, with some showing broad emission lines and others dominated by narrow features (Foley et al., 2016). Of the comparison SNe in Figure 15, SN 2024pxl most closely resembles SN 2014dt, exhibiting narrower lines than the higher-luminosity objects SNe 2020udy and 2008ge. The narrow features seen in the late-phase spectra of Type Iax SNe are thought to originate from the innermost regions (Maeda & Kawabata, 2022; Camacho-Neves et al., 2023). Additional late-time observations of Type Iax SNe across the luminosity distribution will help to understand the origin and nature of these spectral features more precisely.